soybean meal is 14% protein, cornmeal is 7% protein. How many pounds of each should be mixed together in order to get a 280 lb mixture that is 9% protein?
Answer by checkley77(12844)

.14x+.07(280-x)=.09*280
.14x+19.6-.07x=25.2
.07x=25.2-19.6
.07x=5.6
x=5.6/.07
x=80 pounds of soybeans are used.
280-80=200 pounds of cornmeal is used.
Proof:
.14*80+.07*200=.09*280
11.2+14=25.2
25.2=25.2